OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 7-9x-(x+12)<=25? A. x -3 B. x -3 C. -3 x D. A number line with a closed circle on negative 3 and shading to the right.

OpenStudy (anonymous):

Something is wrong with your equation. I don't understand the 25 hanging on the end. Could it be = 25? If so, then 7 - 9x - x - 12 = 25 -10x - 5 = 25 10x = - 30 x = -3

OpenStudy (anonymous):

wait is it x<=-3 or x=>-3?